log3(x-2) = 1+log3(x-4) The logs have a base of 3. How do I solve this?
Answer by scott8148(6628)

log3(x-2) = log3(3) + log3(x-4)

x-2 = 3(x-4)

x-2 = 3x - 12

10 = 2x